# ELA-ZSON: Efficient Layout-Aware Zero-Shot Object Navigation Agent with Hierarchical Planning

### v1

@@ 4.3. LLM-powered Navigation Agent
The LLM directs the operational workflow by selecting the most suitable actions, primarily informed by above mentioned information. Upon entering a new environment, the robot initiates a mapping process, exploring the space frontier-based to gather RGB-D observations and build its scene representation. Once the scene is understood, the agent proceeds with task planning. Continuous monitoring of the trajectory and status ensures that navigation proceeds as intended. If certain conditions arise, such as repeated visits to vertices, excessive time expenditure without reaching the next waypoint, or exceeding re-planning attempts, the LLM either reverts to recalculating the plan or issues an error report. The entire process and primary data flow are shown in Fig.4.
5. Experimental Results
@@
Figure 5. Navigation results on HM3D dataset. The left column shows global planning results, where gray points are entrance vertices in the topological map, red points show the starting positions and yellow points show the destinations. The right columns show local planning results on the ego-centric views.
[…]
The left column shows global planning results, where gray points are entrance vertices in the topological map, red points show the starting positions and yellow points show the destinations. The right columns show local planning results on the ego-centric views.
@@
Figure 6. Real-world experiment results with mobile robotic. The above row shows the navigation success examples. The bottom row shows the unexpected obstacle avoidance examples compared to the built scene memory.
[…]
The above row shows the navigation success examples. The bottom row shows the unexpected obstacle avoidance examples compared to the built scene memory.
@@
Figure 7. Real-world experiment results with Clearpath Jackal. We conducted navigation deployments on different embodied platforms to show its applicability and robustness.
[…]
We conducted navigation deployments on different embodied platforms to show its applicability and robustness.
@@
This section evaluates our proposed ELA-ZSON method through a comparative analysis with SOTA approaches on the widely used MP3D dataset (Ramakrishnan et al., 2021). Additionally, we validate the practicality and robustness of our approach by applying it to real-world multi-room indoor environments.
5.1. Simulation Experiments
@@
Dataset. MP3D (Ramakrishnan et al., 2021) is a high-res photorealistic 3D reconstruction dataset of real-world scenes. We validate our approach on over 20 multi-room indoor scenarios. The dataset is formatted in a Habitat manner and defines six object categories: chairs, couches, potted plants, beds, toilets, and TVs. Notably, our approach considers layout-level information, enabling refined instance navigation experiments by differentiating objects in separate rooms, such as a chair in the living room versus one in the dining room, as distinct instances.
@@
Setups. We employ the Habitat platform as our 3D indoor simulator. Observations resolution from the RGB-D camera is set to $480\times 640$, as used by (Yu et al., 2023). Pose data is provided by an odometry sensor. The automatic exploration process is implemented using a simple policy approach, where the robotic drives continuously along the left frontier while keeping the camera oriented to the right. Frames frequency is set to $15Hz$, following (Yang, 2023). After traversing, the camera is reset to face forward. Our implementation is based on publicly available code examples from the project repository (Puig et al., 2023).
@@
Metrics. We evaluate our approach using Success Rate (SR) and Success Rate weighted by inverse Path Length (SPL), following established metrics in target navigation research (Anderson et al., 2018a). Additionally, we provide instance-level navigation experiments, leveraging our approach’s ability to recognize objects in different rooms as separate instances, capitalizing on the considered layout information.
@@
Comparison Results and Discussions. As shown in Tab.1 and Fig.5, Our ELA-ZSON outperforms existing methods in both SR and SPL. The compared methods include L3MVN (Yu et al., 2023), ESC (Zhou et al., 2023), and VLFM (Yokoyama et al., 2024) which employ LFM-powered mapping of separately frontier, probabilistic commonsense constraints, and visual-language similarity scoring. HOV-SG (Werby et al., 2024) employs global topological map, however, as shown in Fig. 8, the static topological way-points extracted from occupancy cannot ensure efficient path planning or manage unexpected local scene changes which cannot ensure efficiency and robustness. This underscores the advantages of our hierarchical planning strategy, ensuring both effectiveness and efficiency by incorporating refined local waypoints in a global layout context. Moreover, when comparing performances across object and instance navigation tasks, our method excels by setting the final target destination to the appropriate region based on related topology vertices, overcoming limitations of previous methods that ignore layout information.
5.2. Real-world Deployment
@@
Our real-world robotic platforms include two types: one comprises a SLAMTEC mobile base, a Franka Panda arm, and a RealSense D435 camera mounted on the end. The camera is calibrated to the arm’s base coordinates using the easy-hand-eye package. RGB-D images are recorded at $15Hz$, matching the resolution of $480\times 640$ as in (Yu et al., 2023; Yang, 2023). The other one is a Clearpath robotic with the same camera setting. The environment consists of a large-scale, multi-room indoor scene covering approximately $225m^{2}$. It includes a small kitchen tabletop, office area, meeting room, and a hall. As demonstrated in Fig.6, the robot is tasked with navigating to different object instances through long-term paths.
@@
5.3. Ablations
Figure 8. The pre-defined topological voronoi graph is shown on the left, where the green vertices and edges are the actionable path. As shown on the right, Our approach employ sparse layout vertices globally and imperative planning locally rather than simply following a static graph. This leads to efficiency and robustness against scene changes.
[…]
The pre-defined topological voronoi graph is shown on the left, where the green vertices and edges are the actionable path. As shown on the right, Our approach employ sparse layout vertices globally and imperative planning locally rather than simply following a static graph. This leads to efficiency and robustness against scene changes.
Figure 9. Ablations on Global Planning strategy. Results show how the SPL decreases with distance at different settings.
[…]
Results show how the SPL decreases with distance at different settings.
Obstacles w/o Local Planning Ours
Time Cost (s) SR Time Cost (s) SR
@@ 5.3. Ablations
Table 3. Ablations on Local Planning strategy. We evaluate the navigation time cost (only counting the successful navigation) and SR under different numbers of obstacles set in the way after exploration. Each navigation task is evaluated 10 times with similar settings except for the obstacles.
Metrics GPT-3.5 GPT-4o Llama 3 8b DeepSeek-V3
SR 0.812 0.856 0.751 0.867
SPL 0.674 0.797 0.604 0.759
Table 4. Ablations on LLMs the agent employs. It can be seen that different models show different performances, but the metric does not vary too much.
@@
LLM Agent. We ablate different LLMs on the agent role to conduct the entire process. As shown in Tab.4, the SR remains consistent across various LLMs, underscoring the feasibility of employing relatively lightweight models with minimal computational overhead without compromising performance. This indicates that the navigation process can be effectively managed using cost-efficient models without significant reductions.
6. Limitations
@@
This paper introduces ELA-ZSON, an efficient layout-aware zero-shot object navigation approach for multi-room indoor environments, hierarchically planning on both global and local level. Our approach makes progress in addressing globally and locally efficient long-term path planning in complex indoor scenes and ease of deployment without human aid, complex rewards, or costly training. Despite these advancements, our approach currently has two limitations: 1) While local mapping is implemented through imperative learning, it does not fully leverage the potential of the local scene memory for refined obstacle avoidance and improved efficiency. 2) The detected scene changes are not dynamically integrated into the scene representation and topological map. Addressing these limitations by implementing real-time scene updates remains a priority for future work.

### v2

@@ Introduction
The LLM directs the operational workflow by selecting the most suitable actions, primarily informed by above mentioned information. Upon entering a new environment, the robot initiates a mapping process, exploring the space frontier-based to gather RGB-D observations and build its scene representation. Once the scene is understood, the agent proceeds task planning. Continuous monitoring the trajectory and status ensures that navigation proceeds as intended. If errors arise, such as repeated visits, excessive time expenditure without reaching next waypoint, or exceeding re-planning attempts, the LLM either replans or issues an error report. Fig.4 shows entire process and primary data flow.
Experimental Results
@@
Figure 5: Navigation results on MP3D dataset. The left column shows global planning results, where gray points are entrance vertices in the topological map, red points show the starting positions and yellow points show the destinations. The right columns show local planning results on the ego-centric views.
@@
Figure 6: Real-world experiment results with mobile robotic. The above row shows the navigation success examples. The bottom row shows the unexpected obstacle avoidance examples compared to the built scene memory.
@@
Figure 7: Real-world experiment results with Clearpath Jackal. We conducted navigation deployments on different embodied platforms to show its applicability and robustness.
@@
This section evaluates LOG-Nav through a comparative analysis with SOTA approaches on the widely used MP3D dataset (ramakrishnan2021hm3d). Additionally, we validate the practicality and robustness of our approach by applying it to real-world multi-room indoor environments.
Simulation Experiments
@@
Dataset. MP3D (ramakrishnan2021hm3d) is a high-res photorealistic 3D reconstruction dataset of real-world scenes. We validate our approach on over 20 multi-room indoor scenarios. The dataset is formatted in a Habitat manner and defines six object categories: chairs, couches, potted plants, beds, toilets, and TVs. Notably, our approach considers layout-level information, enabling refined instance navigation experiments by differentiating objects in separate rooms, such as a chair in the living room versus one in the dining room, as distinct instances.
@@
Setups. We employ the Habitat platform as our 3D indoor simulator. Observations resolution from the RGB-D camera is set to $480\times 640$, as used by (yu2023l3mvn). Pose data is provided by an odometry sensor. The automatic exploration process is implemented using a simple policy approach, where the robotic drives continuously along the left frontier while keeping the camera oriented to the right. Frames frequency is set to $15Hz$, following (yang2023iplanner). After traversing, the camera is reset to face forward. Our implementation is based on publicly available code examples from the project repository (puig2023habitat3).
@@
Metrics. We evaluate our approach using Success Rate (SR) and Success Rate weighted by inverse Path Length (SPL), following established metrics in target navigation research (anderson2018evaluation). Additionally, we provide instance-level navigation experiments, leveraging our approach’s ability to recognize objects in different rooms as separate instances, capitalizing on the considered layout information.
@@
Comparison Results and Discussions. As shown in Tab.1 and Fig.5, Our Log-Nav outperforms existing methods in both SR and SPL. The compared methods include L3MVN (yu2023l3mvn), ESC (zhou2023esc), and VLFM (yokoyama2024vlfm) which employ LFM-powered mapping of separately frontier, probabilistic commonsense constraints, and visual-language similarity scoring. HOV-SG (werby23hovsg) employs global topological map, however, the static topological way-points extracted from occupancy cannot ensure efficient path planning or manage unexpected local scene changes which cannot ensure efficiency and robustness. This underscores the advantages of our hierarchical planning strategy, ensuring both effectiveness and efficiency by incorporating refined local waypoints in a global layout context. Moreover, when comparing performances across object and instance navigation tasks, our method excels by setting the final target destination to the appropriate region based on related topo-vertices, overcoming limitations of previous methods that ignore layout information.
Further, we compare with zero-shot-object-navigation (ZSON) methods repeatedly within the same workspace to complete object navigation tasks, shown in Tab. 3. All methods shared identical object goal sequences, and we compared the average SPL across different numbers of runs. Our ’explore-then-plan’ approach constructs a complete map only once before performing planning cycles, whereas other ’mapping-during-navigation’ ZSON methods accumulate mapping efforts over repeated executions. Results demonstrate that our mapping cost becomes amortized across planning iterations, particularly beneficial in scenarios with frequent deployments and extensive environments. Furthermore, by leveraging hierarchical environmental priors, our method outperforms incremental mapping approaches.
Methods With Map 1-time Run 10 Runs 20 Runs 30 Runs 60 Runs
ZSON(majumdar2022zson) X 0.102 0.116 0.125 0.119 0.121
PSL(psl) X 0.224 0.215 0.197 0.202 0.199
L3MVN(yu2023l3mvn) $\surd$ 0.186 0.217 0.252 0.281 0.358
ESC(zhou2023esc) $\surd$ 0.249 0.286 0.337 0.351 0.389
LOG-Nav $\surd$ 0.083 0.417 0.533 0.592 0.656
Table 3: SPL comparison with ZSON approaches across different numbers of runs.
Real-world Deployment
@@
Our real-world robotic platforms include two types: one comprises a SLAMTEC mobile base, a Franka Panda arm, and a RealSense D435 camera mounted on the end. The camera is calibrated to the arm’s base coordinates using the easy-hand-eye package. RGB-D images are recorded at $15Hz$, matching the resolution of $480\times 640$ as in (yu2023l3mvn; yang2023iplanner). The other one is a Clearpath robotic with the same camera setting. The environment consists of a large-scale, multi-room indoor scene covering approximately $225m^{2}$. It includes a small kitchen tabletop, office area, meeting room, and a hall. As demonstrated in Fig.6, the robot is tasked with navigating to different object instances through long-term paths.
@@
Ablations
Figure 8: Ablations on Global Planning strategy. Results show how SPL decreases with distance at different settings. ”w/o” means ”without”.
Obstacles w/o Local Planning LOG-Nav (Ours)
Cost (s) SR Cost (s) SR
@@
Table 4: Ablations on Local Planning strategy. We evaluate the navigation time cost (only counting the successful navigation) and SR under different numbers of obstacles set in the way after exploration. Each navigation task is evaluated 10 times with similar settings except for the obstacles.
@@
Limitations
@@
This paper introduces LOG-Nav, an efficient layout-aware object-goal navigation approach for multi-room indoor environments, hierarchically planning on both global and local level. Our approach makes progress in addressing globally and locally efficient long-term path planning in complex indoor scenes and ease of deployment without human aid, complex rewards, or costly training. Despite these advancements, our approach currently has two limitations: 1) While local mapping is implemented through imperative learning, it does not fully leverage the potential of the local scene memory for refined obstacle avoidance and improved efficiency. 2) The detected scene changes are not dynamically integrated into the scene representation and topological map. Addressing these limitations by implementing real-time scene updates remains a priority for future work.
Reproducibility Checklist
Instructions for Authors:
This document outlines key aspects for assessing reproducibility. Please provide your input by editing this .tex file directly.
For each question (that applies), replace the “Type your response here” text with your answer.
Example: If a question appears as
[…]
Type your response here
you would change it to:
[…]
yes
Please make sure to:
  • Replace ONLY the “Type your response here” text and nothing else.
  • Use one of the options listed for that question (e.g., yes, no, partial, or NA).
[…]
1. General Paper Structure
  • $1.$
    Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes/partial/no/NA) yes
  • 1.2.
    Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes/no) yes
  • 1.3.
    Provides well-marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper (yes/no) yes
2. Theoretical Contributions
  • 2.1.
    Does this paper make theoretical contributions? (yes/no) no
    If yes, please address the following points:
    • $2.$
      All assumptions and restrictions are stated clearly and formally (yes/partial/no) Type your response here
    • 2.3.
      All novel claims are stated formally (e.g., in theorem statements) (yes/partial/no) Type your response here
    • 2.4.
      Proofs of all novel claims are included (yes/partial/no) Type your response here
    • 2.5.
      Proof sketches or intuitions are given for complex and/or novel results (yes/partial/no) Type your response here
    • 2.6.
      Appropriate citations to theoretical tools used are given (yes/partial/no) Type your response here
    • 2.7.
      All theoretical claims are demonstrated empirically to hold (yes/partial/no/NA) Type your response here
    • 2.8.
      All experimental code used to eliminate or disprove claims is included (yes/no/NA) Type your response here
3. Dataset Usage
  • 3.1.
    Does this paper rely on one or more datasets? (yes/no) yes
    If yes, please address the following points:
    • 3.2.
      A motivation is given for why the experiments are conducted on the selected datasets (yes/partial/no/NA) yes
    • $3.$
      All novel datasets introduced in this paper are included in a data appendix (yes/partial/no/NA) yes
    • 3.4.
      All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes (yes/partial/no/NA) yes
    • 3.5.
      All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations (yes/no/NA) yes
    • 3.6.
      All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available (yes/partial/no/NA) yes
    • 3.7.
      All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing (yes/partial/no/NA) NA
4. Computational Experiments
  • 4.1.
    Does this paper include computational experiments? (yes/no) yes
    If yes, please address the following points:
    • 4.2.
      This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting (yes/partial/no/NA) yes
    • 4.3.
      Any code required for pre-processing data is included in the appendix (yes/partial/no) yes
    • $4.$
      All source code required for conducting and analyzing the experiments is included in a code appendix (yes/partial/no) partial
    • 4.5.
      All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes (yes/partial/no) yes
    • 4.6.
      All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes/partial/no) yes
    • 4.7.
      If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results (yes/partial/no/NA) yes
    • 4.8.
      This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks (yes/partial/no) yes
    • 4.9.
      This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics (yes/partial/no) yes
    • 4.10.
      This paper states the number of algorithm runs used to compute each reported result (yes/no) yes
    • 4.11.
      Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information (yes/no) yes
    • 4.12.
      The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank) (yes/partial/no) yes
    • 4.13.
      This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments (yes/partial/no/NA) yes